OpenStudy (anonymous):

Simplify each expression. Assume that all variables represent positive real numbers. I'm lost in the issue. Can someone please assist... Thanks in advance

OpenStudy (anonymous):

\[(-x\frac{ 2}{ 3 } \times x\frac{ 1}{ 2 })\div (x \times -x\frac{ 3}{ 2 } )\]

OpenStudy (chihiroasleaf):

can you simplify \(-x \frac{2}{3} \times x \frac{1} {3} \)?

OpenStudy (anonymous):

yes... -x^2 2/3

OpenStudy (anonymous):

i meant 2/9

OpenStudy (chihiroasleaf):

yes ... :) now, what about \[ \Large x \times -x \frac{3}{2} \]

OpenStudy (anonymous):

\[-x ^{2} -1 \frac{ 1 }{ 2} \]

OpenStudy (chihiroasleaf):

are you sure? why is there two terms there? check it again...

OpenStudy (anonymous):

\[-x ^{2}\frac{ 1 }{ 2}\]

OpenStudy (anonymous):

or it may be -2x^2 1/2

OpenStudy (anonymous):

wow i'm messing that all up

OpenStudy (chihiroasleaf):

it's \[\Large x \times -x \frac{3}{2} = -\frac{3}{2} x^{2} \] do you get it?

OpenStudy (anonymous):

but also at the beginning i see that u stated 2/3 and 1/3...it should of been 2/3 and 1/2 ...

OpenStudy (chihiroasleaf):

the first one it's correct \[\Large -\frac{2}{9} x^{2} \]

OpenStudy (chihiroasleaf):

so now you have \[\Large \left( -\frac{2}{9} x^{2} \right) \div \left( -\frac{3}{2} x^2 \right) \] can you simplify?

OpenStudy (anonymous):

yes you can but the original problem was \[(-x\frac{ 2}{ 3 } \times x\frac{ 1}{ 2 })\div (x \times -x\frac{ 3}{ 2 } )\]

OpenStudy (chihiroasleaf):

aahh...., I made mistake it should be 2/6 sorry

OpenStudy (anonymous):

so now i got -1/3 x^2 divided by -3/2 x^2

OpenStudy (chihiroasleaf):

yess... and simplify it ...

OpenStudy (anonymous):

2/9x

OpenStudy (anonymous):

x^2

OpenStudy (anonymous):

or do the x^2 cancel out?

OpenStudy (chihiroasleaf):

what is x^2 divided by x^2

OpenStudy (anonymous):